# Signatu AS, Oslo, Norway

### v1

@@ 4.2 Ontology Engineering Process
The DPVCG consists of experts from multiple disciplines - computer science, law, sociology, authorities, and others. The primary role of its members is to discuss and reach consensus on the scope and information to be represented in DPV. Ontology engineers are then responsible for providing the appropriate modelling of concepts and organisation of DPV as a semantic web resource. While the DPVCG did not formally establish an ontology engineering methodology, practices that were adopted and evolved in the community reflect commonly utilised engineering methodologies of NEON [suarez2011neon] and [poveda2022lot].
@@
Compared to formal ontology engineering methods, this ad-hoc approach lacked the explicit documentation of requirements, competency questions, and use-cases - which would have been expensive to maintain given the limited (regular) participation of volunteers. However, given the stability of the group and continued iteration and adoption of DPV, the group has identified this as an important step to undertake in the future.
The data in spreadsheets was structured to support ontology/taxonomy creation while still being comprehensible to the non-semantic web members. For generating the RDF serialisations, a custom documentation generator was developed which downloaded the spreadsheets, extracted the concepts and serialised them into multiple RDF formats, and generated the corresponding HTMLs. Existing tools such as WIDOCO [garijo2017widoco] were not used due to their limitation in control of outputs - for example DPV has taxonomies of concepts for purposes, personal data, etc. which would be all grouped together in a giant list in WIDOCO outputs, and which in DPV outputs are grouped separately to support adopters see related concepts grouped together. WIDOCO also did not support the dual RDFS+SKOS and OWL modelling outputs of DPV, or the ReSpec template [^9] https://respec.org/docs/ common in W3C outputs.
DPV follows best practices and guidelines established within the semantic web community. Namely, the W3C Best Practices for Publishing Linked Data [^10] Best Practices for Publishing Linked Data (2014) https://www.w3.org/TR/ld-bp/, WIDOCO best practices [garijo2017widoco], W3ID for permanent IRIs, GitHub for version control and collaboration, and Zenodo for archival.
@@ 4.3 Implications of Using SKOS and OWL
As described in the requirements in Section 2, DPV’s application requires supporting semantics where its concepts are used as instances (taxonomy) or as a schema that is instantiated (ontology). Initially, DPV was only provided as an OWL2 ontology. This was expanded upon in DPV v1 which used custom SKOS extensions to define the ’base’ vocabulary with serialisations in RDFS+SKOS and OWL2. In DPV v2, the custom SKOS extensions were removed in favour of RDFS+SKOS as the default serialisation with OWL2 as an alternative serialisation. The RDFS+SKOS serialisation defines concepts as RDFS classes and instances of a top-concept with SKOS used to represent the hierarchy, whereas the OWL2 serialisation uses subclasses to represent the hierarchy.
@@
The resulting implications of this on the adoption and use of DPV are illustrated in the example below. The RDFS+SKOS model is suitable for when DPV is to be used directly in both Tbox and Abox or where the underlying ontology/schema cannot be changed as new concepts are identified in an use-case, whereas the OWL2 model is suitable for when the Tbox and Abox are separate or where there is freedom to change the ontology/schema. Conversely, the OWL2 serialisation is better suited for semantic reasoners than the RDFS+SKOS serialisation based on the underlying formal semantics. By providing both serialisations, DPV enables the adopters to choose the most suitable serialisation that supports their use-case and/or existing implementations, and retains semantic interoperability based on converting between SKOS and OWL2111111Using OWL and SKOS (2008) https://www.w3.org/2006/07/SWD/SKOS/skos-and-owl/master.html.
{minted}
[fontsize=]turtle # Example: A Company is running ’Direct Marketing’ campaigns for # a purpose (CampaignA), within which it runs another special # campaign (CampaignB). This may envolve into another campaign C. # RDFS+SKOS modelling dpv:Marketing a rdfs:Class, dpv:Purpose . dpv:DirectMarketing a dpv:Purpose; skos:broader dpv:Marketing . ex:CampaignA a dpv:Purpose; skos:broader dpv:DirectMarketing . ex:CampaignB a dpv:Purpose; skos:broader ex:CampaignA . # valid statements as all objects are instances of dpv:Purpose ex:Documentation a dpv:Process; dpv:hasPurpose dpv:DirectMarketing .
# OWL2 modelling dpv:Marketing a owl:Class; rdfs:subClassOf dpv:Purpose . dpv:DirectMarketing a owl:Class; rdfs:subClassOf dpv:Marketing . ex:CampaignA a dpv:DirectMarketing . # problem: instance of instance ex:CampaignB a ex:CampaignA . # solution: change CampaignA to subclass ex:CampaignA rdfs:subClassOf dpv:DirectMarketing . # problem: DirectMarketing is not an instance of dpv:Purpose ex:Documentation a dpv:Process; dpv:hasPurpose dpv:DirectMarketing . # solution: blank nodes, but not a good solution ex:Documentation a dpv:Process; dpv:hasPurpose [ a dpv:DirectMarketing ] . # solution: use the instances rather than the class ex:Documentation a dpv:Process; dpv:hasPurpose ex:CampaignA . # problem: does not reflect the terminology used by Company/Law
5 Comparison with Related Work
DPV fills an unique and necessary niche within the state of the art by providing concepts to represent legally relevant information to represent processing of personal data and use of technologies. Here we only describe related works that have a comparable adoption, or are standards, or are outputs of larger projects.
The Open Digital Rights Language (ODRL) [ianella2007open] is a W3C standard for the modelling of policies and agreements which can dictate the permissions, prohibitions, and duties associated with use of data or resources. The ODRL vocabulary also provides specific concepts such as ’consent’ to support commonly used agreements. While there is some overlap between DPV and ODRL, they are complimentary in their objectives and uses. ODRL focuses solely on providing a language for representation of policies and does not focus on specific legal requirements or jurisdictions. DPV’s concepts can thus provide the necessary legal and jurisdictional relevant information within ODRL policies. This has been explored and demonstrated by existing work [esteves_odrl_2021]. The DPVCG intends to establish close collaboration with the ODRL CG through its common members by aligning concepts between DPV and ODRL, and providing guidance for using DPV as an ODRL profile. DPV has also been mentioned by the International Data Spaces Association (IDSA) as a vocabulary of interest to support the implementation of GDPR and regulations in their policies based on ODRL [bader2020international].
LegalRuleML [^12] https://docs.oasis-open.org/legalruleml/legalruleml-core-spec/v1.0/os/legalruleml-core-spec-v1.0-os.html is an OASIS standard providing a ’rule interchange language’ that enables modelling and reasoning tasks based on legal arguments. Similar to ODRL, it focuses on providing a language for expression of ’rules’ and does not provide any taxonomies or model regulations. Similar to ODRL, DPV is also complimentary to LegalRuleML as a vocabulary that can be used to support representation of legal information. There is existing work that has explored the use of LegalRuleML for modelling requirements from the text of GDPR [robaldo2020dapreco], though it has not been maintained nor been extended to support practical requirements related to implementing GDPR.
Gist [^13] https://www.semanticarts.com/gist/ is a “minimalist upper ontology for the enterprise” which provides business concepts with a focus on minimising ambiguity. The Financial Industry Business Ontology [^14] https://spec.edmcouncil.org/fibo/ (FIBO) models concepts relevant in financial business applications, such as contracts and financial transactions. Both Gist and FIBO represent ’industrial ontologies’ and have been developed over a significant portion of time with the involvement of corporate stakeholders. While neither support specific legal requirements or jurisdictions, Gist provides concepts relevant for modelling details about an organisation and FIBO provides modelling of contracts and contract related processes - which DPV does not contain. Work is underway in DPVCG to study the FIBO contract concepts and identify how DPV can support contracts related to personal data processing.
6 Adoption
6.1 Analyses of Citations
In this Section, the adoption of DPVCG’s outputs is evaluated through a citation analysis performed over the DPV v1 publication [panetto_creating_2019]. In this context, 81 publications that cite DPV were found through the Google Scholar service. The gathered results underwent a review and were included in this analysis if deemed pertinent. Duplicated publications, publications without an open-access version and in languages other than English were excluded from this analysis, resulting in 72 publications to be reviewed. Table 1 presents the results of the performed publication evaluation. The publications were evaluated in relation to their use of DPV: publications that reference DPV as a state of the art resource are signalled in the Mention column, that use DPV towards an application or use case in the Use column, and that extend DPV in the Extension column. Publications that contributed their extensions back to DPV (Contributed column) are also marked in Table 1, as well as if the work is applied to a certain domain or sector (Domain column).
Table 1: Citation analysis of academic publications that reference DPV v1 [panetto_creating_2019].
Work Year Mention Use Extension Contributed Domain
[calvaresi_agentbased_2020] 2020 X Health
[lieber_ecodalo_2020] 2020 X Media
[bonatti_machine_2020, matulevicius_method_2020, thalhath_metaprofiles_2020] 2020 X
[calvaresi_personal_2020, krasnashchok_towards_2020, leone_role_2020, ryan_design_2020] 2020 X
[calbimonte_decentralized_2020] 2020 X Health
[debruyne_justtime_2020, pandit_representing_2020, ryan_common_2020] 2020 X X
[esteves_challenges_2021, mcdonald_evaluation_2021, flesch_investigating_2021] 2021 X Health
[grunewald_tira_2021, bonatti_representing_2021, sion_overview_2021, pandit_chapter_2021, hamed_enhancing_2021] 2021 X
[ryan_gdpr_2021, esteves_odrl_2021, ekaputra_semanticenabled_2021, ryan_building_2021, leone_legal_2021] 2021 X
[garcia_towards_2021] 2021 X Smart products
[hickey_gdpr_2021] 2021 X X
[human_data_2022, jesus_consent_2022, human_advanced_2022, sangaroonsilp_mining_2022, pandit_proposals_2022, rasmusen_increasing_2022, esteves_semantifying_2022] 2022 X
[esteves_analysis_2022] 2022 X X
[ryan_support_2022, ryan_dpcat_2022, pandit_semantic_2022, esteves_fostering_2022] 2022 X X
[becker_secondary_2022] 2022 X Health
[esteves_using_2022, debackere_policyoriented_2022, debackere_enforcing_2022, gambarelli_privaframe_2022, kurteva_making_2022] 2022 X
[becher_contra_2022] 2022 X IoT
[hernandez_tikd_2022] 2022 X Health
[pandit_making_2023, bushati_what_2023, breteler_flint_2023, kurteva_relevant_2023, asgarinia_who_2023] 2023 X
[gambarelli_is_2023, grunewald_enabling_2023, bailly_prototyping_2023, tang_helping_2023, esteves_towards_2023, esteves_using_2023, taheri_compliance_2023] 2023 X
[sun_citizencentric_2023, navarrogallinad_evaluating_2023, florea_is_2023, gallinad_usable_2023] 2023 X Health
[kurteva_smashhitcore_2023] 2023 X Smart cities
[gupta_oppo_2023, zichichi_decentralized_2023] 2023 X Media
[esteves_semantics_2023, pandit_towards_2023] 2023 X X
[raza_semantic_2023] 2023 X Health
[pandit_enhancing_2024] 2024 X Health
[herwanto_leveraging_2024] 2024 X
6.2 Projects and Industrial use of DPV
DPV and DPVCG are outputs of the SPECIAL H2020 project and were actively developed and used within the project. In addition to this, DPV was also used in TRAPEZE, MOSAICrOWN, smashHit, FAIRVASC, and PROTECT ITN projects funded under the EU’s H2020 programme which involved both academic and commercial partners. In addition, SPECIAL and TRAPEZE also included a Data Protection Authority who provided legal expertise in implementation and design of ontologies. Of these, TRAPEZE actively contributed back to the DPV and was instrumental in identifying the design structure where both RDFS+SKOS and OWL serialisations were developed, and to assist in developing a framework to provide DPV concepts with multilingual documentation.
In addition to the use of DPV in industrial context in above projects, companies that actively utilise DPV include Signatu – which develops legal compliance solutions, and Inrupt – which develops the Solid specifications and implementations. The DPVCG has received contributions from both, with Signatu being an active and regular contributor and providing legal expertise and practical requirements for industrial applications.
6.3 Use of DPV in Standards
ISO/IEC 27560:2023 is a Technical Specification (TS) describing a ’consent record information structure’, which defines the specific information to be maintained in consent records. DPV’s consent modelling played a significant part in the development of this standard based on sharing the knowledge [^15] Harshvardhan J. Pandit was an active contributor to the development of ISO/IEC 27560:2023 through the National Standards Authority of Ireland (NSAI) of legal and semantic requirements regarding consent records and receipts. The annex of 27560:2023 provides examples of consent records and receipts using JSON-LD - where DPV is explicitly referenced in the document and its concepts are used as to define the schema (e.g. hasPurpose and instances (e.g. Service Provision).
The IEEE P7012 Working Group is developing a specification to define how “personal privacy terms are proffered and how they can be read and agreed to by machines”. It explicitly references DPV as the vocabulary to describe activities regarding processing of personal data in an interoperability manner.
7 Planned Future Developments
The DPV’s status as a state of the art resource depends on its ability to be up to date with regulatory developments as well as the evolving landscape of technologies - especially those that relate to AI. In addition, its adoption relies on the ability for stakeholders to understand its value and have convenient information on how to integrate it within different kinds of projects. To support this, the DPVCG has identified the necessity to provide a rich set of examples that illustrate not only the concepts but also the semantic and legal implications of using them, and developing guides and tutorials for specific applications. To assist in refining the ontology engineering process, the group is implementing processes to document requirements and use-cases in a more formal manner as they are discussed and resolved.
7.1 Extending DPV
The DPVCG plans to refine its extension providing taxonomies of technology related concepts such as how it was provisioned, its network connectivity requirements, and to model practical necessities such as whether it is a ’cloud service’ or ’runs only on device’. In addition, recent discussions have also identified the necessity to specifically model AI related concepts as a further extension of the existing Technology extension to provide AI specific taxonomies, for example - AI techniques, capabilities, lifecycle stages, risks and risk sources, and to enable stakeholders to express complex concepts such as generative AI in a manner that supports regulatory requirements and is based on recent developments of ISO standards. The DPVCG is also working on providing a risk assessment and risk management vocabulary based on the ISO 31000 series to support the various risk related processes.
The DPVCG is also continuing its efforts to develop vocabularies to represent key ’data and AI regulations’ notably in EU the Digital Services Act (DSA), Data Markets Act (DMA), Data Act, Network Information Security Directive (NIS2), and the recently finalised AI Act. In addition, the DPVCG is keen to demonstrate the pan-jurisdictional capabilities of DPV by modelling regulations in specific countries such as Ireland, USA, and UK.
To support the application of DPV in regulatory environments, the DPVCG is developing guides based on existing work that utilises DPV to support GDPR implementations. These include Records of Processing Activities (ROPA) [ryan_dpcat_2022], Data Protection Impact Assessment (DPIA) [pandit_semantic_2022], Data Breaches [pandit_towards_2023], and Rights management. To support the implementation of decentralised and user-centric applications of DPV, such as those envisioned in the IDSA, IEEE P7012 and Solid, the DPVCG is developing vocabularies and guides - for example implementation of ISO Standards for (semantic) privacy notices. In addition, the DPV is also looking to incorporate the Standard Data Protection Model [^16] https://www.bfdi.bund.de/EN/Fachthemen/Inhalte/Technik/SDM.html provided by the German data protection authority regarding implementation of concrete technical and organisational measures.
7.2 Tooling Support for Validation and Reasoning
The focus of DPV so far has been on enabling information representation. To ensure that this information is ’correct’ and ’complete’ the group is also exploring defining specifications which guide how information should be represented, and how it could be verified - for example through SHACL shapes. While this is not a sufficient replacement for formal semantics or the creation of a ’language’, it is intended to encourage development of common tooling and infrastructure through which barriers can be reduced to utilise DPV and semantic technologies regarding regulatory processes.
8 Conclusion
@@
DPV version 2 represents a significant milestone in the development of a vocabulary to support legally relevant processes, and an expansion of scope to include multiple regulations and jurisdictions, as well as recent advances in AI and data sharing architectures such as Solid and IDSA. The DPVCG, in continuing to develop the DPV, welcomes more participation and contributions to support its vision of providing an interoperable vocabulary that provides value and supports making legal compliance processes more efficient and aligned for all stakeholders.
@@
The DPVCG was established as part of the SPECIAL H2020 Project, which received funding from the European Union’s Horizon 2020 research and innovation programme under Grant#731601 from 2017 to 2019. Harshvardhan J. Pandit was funded to work on DPV from 2020 to 2022 by the Irish Research Council’s Government of Ireland Postdoctoral Fellowship Grant#GOIPD/2020/790. The ADAPT SFI Centre for Digital Media Technology is funded by Science Foundation Ireland through the SFI Research Centres Programme and is co-funded under the European Regional Development Fund (ERDF) through Grant#13/RC/2106 (2018 to 2020) and Grant#13/RC/2106_P2 (2021 onwards). The contributions of Piero Bonatti and Luigi Sauro to the DPVCG have been funded by the European Union’s Horizon 2020 research and innovation programme under grant agreement N. 731601 (project SPECIAL) until 2019, and under Grant#883464 (project TRAPEZE) from 2020 until 2023. The contributions of Beatriz Esteves and Delaram Golpayegani have received funding through the PROTECT ITN Project from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant#813497.

### v2

@@ 6.2 Ontology Engineering Process
The DPVCG consists of experts from multiple disciplines — computer science, law, sociology, authorities, and others. The primary role of its members is to discuss and reach consensus on the scope and information to be represented in DPV. Ontology engineers are then responsible for providing the appropriate modelling of concepts and organisation of DPV as a semantic web resource. While the DPVCG did not formally establish an ontology engineering methodology, practices that were adopted and evolved in the community reflect commonly utilised engineering methodologies of NeOn [suarez2011neon] and LOT [poveda2022lot].
@@
Compared to formal ontology engineering methods, this ad-hoc approach lacked the explicit documentation of requirements, competency questions, and use-cases — which would have been expensive to maintain given the limited (regular) participation of volunteers. However, given the stability of the group and continued iteration and adoption of DPV, the group has identified this as an important step to undertake in the future.
The data in spreadsheets was structured to support ontology/taxonomy creation while still being comprehensible to the non-semantic web members. For generating the RDF serialisations, a custom documentation generator was developed which downloaded the spreadsheets, extracted the concepts and serialised them into multiple RDF formats, and generated the corresponding HTMLs. Existing tools such as WIDOCO [garijo2017widoco] were not used due to their limitation in control of outputs — for example DPV has taxonomies of concepts for purposes, personal data, etc. which would be all grouped together in a giant list in WIDOCO outputs, and which in DPV outputs are grouped separately to support adopters seeing related concepts grouped together. WIDOCO also did not support the dual RDFS+SKOS and OWL modelling outputs of DPV, or the ReSpec template [^20] https://respec.org/docs/ common in W3C outputs.
DPV follows best practices and guidelines established within the semantic web community. Namely, the W3C Best Practices for Publishing Linked Data [^21] Best Practices for Publishing Linked Data (2014) https://www.w3.org/TR/ld-bp/, WIDOCO best practices [garijo2017widoco] with OOPS! [^22] https://oops.linkeddata.es/ and FOOPS! [^23] https://foops.linkeddata.es/ — DPV’s score has false negatives regarding use of SKOS annotations, this has been reported to the maintainers for evaluation, W3ID [^24] https://w3id.org/ for permanent IRIs, GitHub [^25] https://github.com/w3c/dpv for version control and collaboration, and Zenodo [^26] https://doi.org/10.5281/zenodo.12505840 for archival.
@@ 6.3 Implications of Using SKOS and OWL
As described in the requirements in Section 2, use-cases utilising DPV involve cases where its concepts are used as instances (taxonomy) or as a schema that is instantiated (ontology). Initially, DPV was only provided as an OWL2 ontology. This was expanded upon in DPV 1.0 which used custom SKOS extensions to define the ‘base’ vocabulary with serialisations in RDFS+SKOS and OWL2 with the goal of supporting both categories of use-cases. In DPV 2.0, the custom SKOS extension was removed in favour of using RDFS+SKOS as standards for the default serialisation and providing an alternative serialisation for OWL2.
The RDFS+SKOS serialisation defines concepts as instances of rdfs:Class and skos:Concept. To create a hierarchical taxonomy, the concepts are represented as instances of a top-concept (e.g., dpv:Marketing as an instance of dpv:Purpose) and skos:broader/narrower is used to define the relation between instances. In OWL2, concepts are defined as owl:Class and the hierarchy is defined using rdfs:subClassOf. The justification for why DPV is provided with two different semantics is illustrated in Figure 2.
Figure 2: Semantic implications of using DPV with RDFS+SKOS and OWL semantics
The Figure compares the implementation of RDFS+SKOS and OWL2 for an use-case that uses the purpose taxonomy from DPV. The use-case involves three steps of documentation (not shown in the figure) where the organisation first records that its planned purpose is ‘Direct Marketing’, and then later it creates ‘CampaignA’ as a specific form of direct marketing, and even later creates ‘CampaignB’ as a specific part of ‘CampaignA’ direct marketing.
In RDFS+SKOS, both dpv:DirectMarketing and ex:CampaignA are both defined as an instance of dpv:Purpose, and associated with each other using skos:broader. Using these as the object with property dpv:hasPurpose is correct as its range is (instances of) dpv:Purpose. Later, when ex:CampaignB is introduced, it does not require changes to DPV or the use-case graph (e.g., to convert instances into classes) as the new concept is also defined as an instance of dpv:Purpose and can be used with the property.
In OWL2, all DPV concepts are defined as instances of owl:Class and associated with each other using rdfs:subClassOf. The use-case concepts can now be declared as either subclasses of DPV concepts or as instances. In either case, DPV concepts cannot be directly used with the property dpv:hasPurpose as they are not instances of dpv:Purpose (shown in red). If the use-case creates instances — as shown where ex:CampaignA is created as an instance of dpv:DirectMarketing — then the use-case concept can be correctly used with dpv:hasPurpose. However, when ex:CampaignB is later introduced, it cannot be subclass of ex:CampaignA as ‘subclass of instance’ is undefined [^27] It might be tempting to use punning here — but it would not be correct. Punning enables use of the same concept as class and instance in separate contexts whereas we have a concept as class and instance in the same context — which is undefined. in OWL2. The relation between the two use-case concepts thus must be defined using either SKOS (thereby mixing SKOS and OWL2) or by using or creating another property, e.g., dct:hasPart. At this point, the use-case should reengineer its ontology by creating a class representing ex:CampaignA and creating the necessary subclass and instances to represent the relationship with ex:CampaignB. This however does not solve the issue with directly using DPV concepts as instances.
@@
Thus, the RDFS+SKOS model is suitable for when DPV is to be used as a controlled taxonomy with a ‘lightweight ontology’ that supports extending the taxonomy in use-cases. Conversely, the OWL2 model is better suited for cases where formal reasoning is needed, and where sufficient ontology engineering capabilities exist to address changes in use-cases. By providing both serialisations, DPV enables the adopters to choose the most suitable serialisation that supports their use-case and/or existing implementations, and retains semantic interoperability based on converting between SKOS and OWL2282828Using OWL and SKOS (2008) https://www.w3.org/2006/07/SWD/SKOS/skos-and-owl/master.html.
7 Conclusion & Future Work
@@
DPV version 2 represents a significant milestone in the development of a vocabulary to support legally relevant processes across multiple regulations and jurisdictions, as well as recent advances in AI and data sharing regulations such as through EU’s DGA and AI Act, and in architectures such as Solid and IDSA. The DPVCG, in continuing to develop the DPV, welcomes more participation and contributions to support its vision of providing an interoperable vocabulary that provides value and supports making legal compliance processes more efficient and aligned for all stakeholders.
The DPVCG plans to refine its TECH and AI extensions based on existing works [golpayeganiAIROOntologyRepresenting2022, golpayeganiAICardsApplied2024] providing taxonomies for AI techniques, capabilities, lifecycle stages, risks and risk sources, and to enable stakeholders to express specific use-cases (e.g., involving generative AI) in a manner that supports requirements for EU AI Act and ISO standards [golpayeganiAICardsApplied2024]. The DPVCG is also continuing its efforts to develop vocabularies to represent key ‘data and AI regulations’ notably in EU the Digital Services Act (DSA), Data Markets Act (DMA), Data Act, and Data Spaces, and modelling laws in other jurisdictions, e.g., Ireland, USA, and UK.
To support the application of DPV in regulatory environments, the DPVCG is developing guides based on existing work that utilises DPV to support GDPR implementations. These include Records of Processing Activities (ROPA) [ryan_dpcat_2022], Data Protection Impact Assessment (DPIA) [pandit_semantic_2022], Data Breaches [pandit_towards_2023], and Rights management. To support the implementation of decentralised and user-centric applications of DPV, such as those envisioned in the IDSA, IEEE P7012 and Solid, the DPVCG is developing vocabularies and guides — for example implementation of ISO Standards for (semantic) privacy notices. In addition, the DPV is also looking to incorporate the Standard Data Protection Model [^29] https://www.bfdi.bund.de/EN/Fachthemen/Inhalte/Technik/SDM.html provided by the German data protection authority regarding implementation of concrete technical and organisational measures.
Resource Availability Statement:
The source and releases for the DPV are available via GitHub: https://github.com/w3c/dpv and have been deposited in Zenodo for long-term archival: https://doi.org/10.5281/zenodo.12505840.
@@
The DPVCG was established as part of the SPECIAL H2020 Project, which received funding from the European Union’s Horizon 2020 research and innovation programme under Grant#731601 from 2017 to 2019. Harshvardhan J. Pandit was funded to work on DPV from 2020 to 2022 by the Irish Research Council’s Government of Ireland Postdoctoral Fellowship Grant#GOIPD/2020/790. The ADAPT SFI Centre for Digital Media Technology is funded by Science Foundation Ireland through the SFI Research Centres Programme and is co-funded under the European Regional Development Fund (ERDF) through Grant#13/RC/2106 (2018 to 2020) and Grant#13/RC/2106_P2 (2021 onwards). The contributions of Piero Bonatti and Luigi Sauro to the DPVCG have been funded by the European Union’s Horizon 2020 research and innovation programme under grant agreement N. 731601 (SPECIAL) until 2019, and under Grant#883464 (TRAPEZE) from 2020 until 2023. The contributions of Beatriz Esteves and Delaram Golpayegani have received funding through the PROTECT ITN Project from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant#813497. Beatriz is also funded by SolidLab Vlaanderen (Flemish Government, EWI and RRF project VV023/10). The contributions of Julian Flake have received funding from the German Federal Ministry of Education and Research (BMBF) grant#16KIS1298 (AI-NET PROTECT), from the European Union’s Horizon Europe Framework Programme grant#101129822 (TITAN) and from the European Union’s Digital Europe Programme grant#101123471 (EDGE-Skills).